OpenStudy (anonymous):

Will fan and medal!!! Graph as a transformation of the parent function: y = -2)^x+3 What is the parent function? how do I graph it?

Directrix (directrix):

Is this ---> y = -2)^x+3 supposed to be y = (-2)^x + 3 ? Or, maybe y = -2^x+3

OpenStudy (anonymous):

no... sorry y = -2(5)^x

Directrix (directrix):

The negative portion of the (-2) will cause the graph to do a flip and move from quads I and II to quads III and IV. It is a reflection about the x-axis with the proviso that the functional values of y = -2(5)^x are two times as small as the corresponding y values of the original parent function y = -(5)^x. You can see the two graphs here: http://www.wolframalpha.com/input/?i=+y+%3D%285%29%5Ex+%2C++y+%3D+-2%285%29%5Ex @K_V8